# Double robust estimation of functional outcomes with data missing at random

### v2

@@ Appendix A Asymptotic results for finite-dimensional distributions
$\mathbb{E}\left(\frac{\partial\psi_{i}(\boldsymbol{\theta})}{\partial\boldsymbol{\theta}}\biggr{\rvert}_{\boldsymbol{\theta}_{0}}\right)=\begin{bmatrix}-\textbf{I}_{T\times T}&\textbf{0}_{T\times Tp}&\textbf{0}_{T\times p}\\ \textbf{0}_{Tp\times T}&\textbf{I}_{T\times T}\otimes\boldsymbol{\Pi}_{p\times p}&\textbf{0}_{p\times p}\\ \textbf{0}_{p\times T}&\textbf{0}_{p\times Tp}&\boldsymbol{\Phi}_{p\times p}\end{bmatrix}$,
where $\boldsymbol{\Pi}=\mathbb{E}\left[Z_{i}\textbf{x}_{i}\textbf{x}_{i}^{\top}\right]$ and $\boldsymbol{\Phi}=\mathbb{E}\left[\tau(\textbf{x}_{i}^{\top}\boldsymbol{\gamma}_{0})(1-\tau(\textbf{x}_{i}^{\top}\boldsymbol{\gamma}_{0}))\textbf{x}_{i}\textbf{x}_{i}^{\top}\right]$. To see why this is the case, note that if the missingness model (2) is correctly specified, the second block in the first row of the Jacobian simplifies as follows:
$\mathbb{E}\left(\frac{\tau(\textbf{x}_{i}^{\top}\boldsymbol{\gamma}_{0})-Z_{i}}{\tau(\textbf{x}_{i}^{\top}\boldsymbol{\gamma}_{0})}(\textbf{I}_{T\times T}\otimes\textbf{x}_{i}^{\top})\right)$
$=\mathbb{E}_{\textbf{x}}\left(\mathbb{E}\left(\frac{\tau(\textbf{x}_{i}^{\top}\boldsymbol{\gamma}_{0})-Z_{i}}{\tau(\textbf{x}_{i}^{\top}\boldsymbol{\gamma}_{0})}(\textbf{I}_{T\times T}\otimes\textbf{x}_{i}^{\top})\biggr{\rvert}\textbf{x}_{i}\right)\right)=0$
If the outcome model (A.1) is correctly specified, the third block in the first row of the Jacobian simplifies as follows:
$\mathbb{E}\left(-Z_{i}e^{-\textbf{x}_{i}^{\top}\boldsymbol{\gamma}_{0}}\boldsymbol{\epsilon}_{i}\textbf{x}_{i}^{\top}\right)=\mathbb{E}\boldsymbol{\epsilon}_{i}\mathbb{E}\left(-Z_{i}e^{-\textbf{x}_{i}^{\top}\boldsymbol{\gamma}_{0}}\textbf{x}_{i}^{\top}\right)=\textbf{0}.\\$
Next, calculate $\mathbb{E}[\psi_{i}(\boldsymbol{\theta}_{0})\psi_{i}(\boldsymbol{\theta}_{0})^{\top}]$. Still assuming both working models are correctly misspecified, we have:
$\mathbb{E}[\psi_{1,i}(\boldsymbol{\theta}_{0})\psi_{1,i}(\boldsymbol{\theta}_{0})^{\top}]$
$=\mathbb{C}\text{ov}\left[(\textbf{I}_{T\times T}\otimes\textbf{x}_{i}^{\top})\emph{vec}(\textbf{B}_{0})\right]+\mathbb{E}\left[\frac{Z_{i}}{\tau^{2}(\textbf{x}_{i}^{\top}\boldsymbol{\gamma}_{0})}\boldsymbol{\epsilon}_{i}\boldsymbol{\epsilon}_{i}^{\top}\right]$
$=\textbf{B}^{\top}_{0}\boldsymbol{\Sigma}_{x}\textbf{B}_{0}+\mathbb{E}[\tau^{-1}(\textbf{x}_{i}^{\top}\boldsymbol{\gamma}_{0})]\boldsymbol{\Sigma}_{\epsilon},$
@@
$\mathbb{E}\left[\frac{1}{\tau(\textbf{x}_{i}^{\top}\boldsymbol{\gamma}_{0})}\right]=\frac{1}{\pi}\left(1-\frac{\mathbb{V}(\tau(\textbf{x}_{i}^{\top}\boldsymbol{\gamma}_{0}))}{[\mathbb{E}\tau(\textbf{x}_{i}^{\top}\boldsymbol{\gamma}_{0})]^{2}}\right)=\frac{1}{\pi}\left(1-\frac{\mathbb{V}(\tau(\textbf{x}_{i}^{\top}\boldsymbol{\gamma}_{0}))}{\pi^{2}}\right),$
@@
$\mathbb{V}(\tau(\textbf{x}_{i}^{\top}\boldsymbol{\gamma}_{0}))$ $=\mathbb{E}_{\textbf{x}}[\tau(\textbf{x}_{i}^{\top}\boldsymbol{\gamma}_{0})^{2}]-\pi^{2}.$
@@
$\mathbb{E}[\psi_{1,i}(\boldsymbol{\theta}_{0})\psi_{2,i}(\boldsymbol{\theta}_{0})^{\top}]$
$=\mathbb{E}\left[\left((\textbf{I}_{T\times T}\otimes\textbf{x}_{i}^{\top})\emph{vec}(\textbf{B}_{0})-\boldsymbol{\mu}_{y}^{0}+\frac{Z_{i}}{\tau(\textbf{x}_{i}^{\top}\boldsymbol{\gamma}_{0})}\boldsymbol{\epsilon}_{i}\right)\left(\boldsymbol{\epsilon}_{i}\otimes Z_{i}\textbf{x}_{i}\right)^{\top}\right]$
$=\mathbb{E}\left[\frac{Z_{i}}{\tau(\textbf{x}_{i}^{\top}\boldsymbol{\gamma}_{0})}\boldsymbol{\epsilon}_{i}\left(\boldsymbol{\epsilon}_{i}\otimes Z_{i}\textbf{x}_{i}\right)^{\top}\right]=\mathbb{E}\left[\frac{Z_{i}}{\tau(\textbf{x}_{i}^{\top}\boldsymbol{\gamma}_{0})}(\boldsymbol{\epsilon}_{i}\boldsymbol{\epsilon}_{i}^{\top}\otimes\textbf{x}_{i}^{\top})\right]$
$=\boldsymbol{\Sigma}_{\epsilon}\otimes\boldsymbol{\mu}_{x}^{\top}$
@@
$\mathbb{E}[\psi_{1,i}(\boldsymbol{\theta}_{0})\psi_{3,i}(\boldsymbol{\theta}_{0})^{\top}]$
$=\mathbb{E}\left[\left((\textbf{I}_{T\times T}\otimes\textbf{x}_{i}^{\top})\emph{vec}(\textbf{B}_{0})-\boldsymbol{\mu}_{y}^{0}+\frac{Z_{i}}{\tau(\textbf{x}_{i}^{\top}\boldsymbol{\gamma})}\boldsymbol{\epsilon}_{i}\right)(Z_{i}-\tau(\textbf{x}_{i}^{\top}\boldsymbol{\gamma}))\textbf{x}_{i}^{\top}\right]=\textbf{0}.$
@@
$\mathbb{E}[\psi_{2,i}(\boldsymbol{\theta}_{0})\psi_{2,i}(\boldsymbol{\theta}_{0})^{\top}]$
$=\mathbb{E}\left[(\boldsymbol{\epsilon}_{i}\otimes Z_{i}\textbf{x}_{i})(\boldsymbol{\epsilon}_{i}\otimes Z_{i}\textbf{x}_{i})^{\top}\right]=\mathbb{E}\boldsymbol{\epsilon}_{i}\boldsymbol{\epsilon}_{i}^{\top}\otimes\left[Z_{i}\textbf{x}_{i}\textbf{x}_{i}^{\top}\right]$
$=\mathbb{E}\left[\boldsymbol{\epsilon}_{i}\boldsymbol{\epsilon}_{i}^{\top}\right]\otimes\mathbb{E}\left[Z_{i}\textbf{x}_{i}\textbf{x}_{i}^{\top}\right]=\boldsymbol{\Sigma}_{\epsilon}\otimes\boldsymbol{\Pi}$
@@
$\mathbb{E}[\psi_{2,i}(\boldsymbol{\theta}_{0})\psi_{3,i}(\boldsymbol{\theta}_{0})^{\top}]=\mathbb{E}\left[(\boldsymbol{\epsilon}_{i}\otimes Z_{i}\textbf{x}_{i})(Z_{i}-\tau(\textbf{x}_{i}^{\top}\boldsymbol{\gamma}))\textbf{x}_{i}^{\top}\right]=\textbf{0}.$
Since $\hat{\boldsymbol{\gamma}}$ is a maximum likelihood estimator, the blocks in $\mathbb{E}[\psi_{3,i}(\boldsymbol{\theta}_{0})\psi_{3,i}(\boldsymbol{\theta}_{0})^{\top}]$ are identical to the expected value of the corresponding blocks in the Jacobian. According to the M-estimation approach, we have $\sqrt{n}(\hat{\boldsymbol{\theta}}-\boldsymbol{\theta})\overset{d}{\rightarrow}\mathcal{N}(\textbf{0},\boldsymbol{\Sigma_{\theta}})$, where, by the sandwich estimator formula, the asymptotic variance of $\hat{\boldsymbol{\theta}}$ is
$\mathbb{E}\left(\frac{\partial\psi_{i}(\boldsymbol{\theta})}{\partial\boldsymbol{\theta}}\biggr{\rvert}_{\boldsymbol{\theta}_{0}}\right)^{-1}\mathbb{E}[\psi_{i}(\boldsymbol{\theta}_{0})\psi_{i}(\boldsymbol{\theta}_{0})^{\top}]\left\{\mathbb{E}\left(\frac{\partial\psi_{i}(\boldsymbol{\theta})}{\partial\boldsymbol{\theta}}\biggr{\rvert}_{\boldsymbol{\theta}_{0}}\right)^{-1}\right\}^{\top}$
$=\begin{bmatrix}\textbf{B}^{\top}_{0}\boldsymbol{\Sigma}_{x}\textbf{B}_{0}+\mathbb{E}[\tau^{-1}(\textbf{x}_{i}^{\top}\boldsymbol{\gamma}_{0})]\boldsymbol{\Sigma}_{\epsilon}&-\boldsymbol{\Sigma}_{\epsilon}\otimes\boldsymbol{\mu}_{x}^{\top}\boldsymbol{\Pi}^{-1}&\textbf{0}\\ \boldsymbol{\Sigma}_{\epsilon}\otimes\boldsymbol{\Pi}^{-1}\boldsymbol{\mu}_{x}&\boldsymbol{\Sigma}_{\epsilon}\otimes\boldsymbol{\Pi}^{-1}&\textbf{0}\\ \textbf{0}&\textbf{0}&\boldsymbol{\Phi}^{-1}\end{bmatrix}$.
@@
Lemma A.2.
Assume the working model (1) is correctly specified, i.e., $E(\textbf{y}_{i}\mid\textbf{x}_{i})=\textbf{B}^{\top}\textbf{x}_{i}$. Then, the OR estimator (A.2) has the following asymptotic multivariate normal distribution:
$\sqrt{n}(\hat{\boldsymbol{\mu}}_{OR}-\boldsymbol{\mu}_{y})\overset{d}{\rightarrow}\mathcal{N}\left(\textbf{0},\textbf{B}^{\top}\boldsymbol{\Sigma}_{x}\textbf{B}+\boldsymbol{\mu}_{x}^{\top}\boldsymbol{\Pi}^{-1}\boldsymbol{\mu}_{x}\boldsymbol{\Sigma}_{\epsilon}\right),$
where $\boldsymbol{\Pi}=\mathbb{E}\left[Z_{i}\textbf{x}_{i}\textbf{x}_{i}^{\top}\right]$.
@@
Here, we apply the same approach to find the limiting distribution of outcome regression imputed estimator. The outcome regression imputed estimator is defined as $\hat{\boldsymbol{\mu}}_{OR}=n^{-1}\sum_{i=1}^{n}\hat{\textbf{B}}^{\top}\textbf{x}_{i}$. Let $\boldsymbol{\theta}=(\boldsymbol{\mu}_{y},\emph{vec}(\textbf{B}))^{\top}$. Define equations for calculating the estimator as:
$\psi_{1,i}(\boldsymbol{\theta})=(\textbf{I}_{T\times T}\otimes\textbf{x}_{i}^{\top})\emph{vec}(\textbf{B})-\boldsymbol{\mu}_{y},$ (A.7)
$\psi_{2,i}(\boldsymbol{\theta})=(\textbf{y}_{i}-(\textbf{I}_{T\times T}\otimes\textbf{x}_{i}^{\top})\emph{vec}(\textbf{B}))\otimes Z_{i}\textbf{x}_{i}.$ (A.8)
Calculate the Jacobian of $\psi_{i}(\boldsymbol{\theta})$ and evaluate it at the true value $\boldsymbol{\theta}_{0}$. The first row of Jacobian is
$\frac{\partial\psi_{1,i}(\boldsymbol{\theta})}{\partial\boldsymbol{\theta}^{\top}}\biggr{\rvert}_{\boldsymbol{\theta}_{0}}=\left[-\textbf{I}_{T\times T},\textbf{I}_{T\times T}\otimes\textbf{x}_{i}^{\top}\right].$
@@
$\frac{\partial\psi_{2,i}(\boldsymbol{\theta})}{\partial\boldsymbol{\theta}^{\top}}\biggr{\rvert}_{\boldsymbol{\theta}_{0}}=\left[\textbf{0}_{Tp\times T},-\textbf{I}_{T\times T}\otimes Z_{i}\textbf{x}_{i}\textbf{x}_{i}^{\top}\right].$
@@
$\mathbb{E}\left(\frac{\partial\psi_{i}(\boldsymbol{\theta})}{\partial\boldsymbol{\theta}}\biggr{\rvert}_{\boldsymbol{\theta}_{0}}\right)=\begin{bmatrix}-\textbf{I}_{T\times T}&\textbf{I}_{T\times T}\otimes\boldsymbol{\mu}_{x}^{\top}\\ \textbf{0}_{Tp\times T}&-\textbf{I}_{T\times T}\otimes\boldsymbol{\Pi}_{p\times p}\end{bmatrix}$,
where $\boldsymbol{\Pi}=\mathbb{E}\left[Z_{i}\textbf{x}_{i}\textbf{x}_{i}^{\top}\right]$. By the M-estimation approach, the asymptotic variance of $\hat{\boldsymbol{\theta}}$ is
$\mathbb{E}\left(\frac{\partial\psi_{i}(\boldsymbol{\theta})}{\partial\boldsymbol{\theta}}\biggr{\rvert}_{\boldsymbol{\theta}_{0}}\right)^{-1}\mathbb{E}[\psi_{i}(\boldsymbol{\theta}_{0})\psi_{i}(\boldsymbol{\theta}_{0})^{\top}]\left\{\mathbb{E}\left(\frac{\partial\psi_{i}(\boldsymbol{\theta})}{\partial\boldsymbol{\theta}}\biggr{\rvert}_{\boldsymbol{\theta}_{0}}\right)^{-1}\right\}^{\top}$
$=\begin{bmatrix}\textbf{B}^{\top}_{0}\boldsymbol{\Sigma}_{x}\textbf{B}_{0}+\boldsymbol{\mu}_{x}^{\top}\boldsymbol{\Pi}^{-1}\boldsymbol{\mu}_{x}\boldsymbol{\Pi}_{\epsilon}&\boldsymbol{\Sigma}_{\epsilon}\otimes\boldsymbol{\mu}^{\top}\boldsymbol{\Pi}^{-1}\\ \boldsymbol{\Sigma}_{\epsilon}\otimes\boldsymbol{\Pi}^{-1}\boldsymbol{\mu}_{x}&\boldsymbol{\Sigma}_{\epsilon}\otimes\boldsymbol{\Pi}^{-1}\end{bmatrix}$.
@@
Lemma A.1 states the double robustness property of the DR estimator (e.g. Tsiatis, 2006). However, up to our knowledge, the covariance matrices of the pointwise OR- and DR-estimators in the context of multivariate outcomes do not appear in the literature elsewhere. These covariance matrices are necessary to provide simultaneous inference as presented in Section 2.3.

### v3

@@ Appendix A Asymptotic results for finite-dimensional distributions
$\mathbb{E}\left(\frac{\partial\psi_{i}(\bm{\theta})}{\partial\bm{\theta}}\biggr\rvert_{\bm{\theta}_{0}}\right)=\begin{bmatrix}-\textbf{I}_{T\times T}&\textbf{0}_{T\times Tp}&\textbf{0}_{T\times p}\\ \textbf{0}_{Tp\times T}&\textbf{I}_{T\times T}\otimes\bm{\Pi}_{p\times p}&\textbf{0}_{p\times p}\\ \textbf{0}_{p\times T}&\textbf{0}_{p\times Tp}&\bm{\Phi}_{p\times p}\end{bmatrix}$,
where $\bm{\Pi}=\mathbb{E}\left[Z_{i}\textbf{x}_{i}\textbf{x}_{i}^{\top}\right]$ and $\bm{\Phi}=\mathbb{E}\left[\tau(\textbf{x}_{i}^{\top}\bm{\gamma}_{0})(1-\tau(\textbf{x}_{i}^{\top}\bm{\gamma}_{0}))\textbf{x}_{i}\textbf{x}_{i}^{\top}\right]$. To see why this is the case, note that if the missingness model (2) is correctly specified, the second block in the first row of the Jacobian simplifies as follows:
$\mathbb{E}\left(\frac{\tau(\textbf{x}_{i}^{\top}\bm{\gamma}_{0})-Z_{i}}{\tau(\textbf{x}_{i}^{\top}\bm{\gamma}_{0})}(\textbf{I}_{T\times T}\otimes\textbf{x}_{i}^{\top})\right)$
$=\mathbb{E}_{\textbf{x}}\left(\mathbb{E}\left(\frac{\tau(\textbf{x}_{i}^{\top}\bm{\gamma}_{0})-Z_{i}}{\tau(\textbf{x}_{i}^{\top}\bm{\gamma}_{0})}(\textbf{I}_{T\times T}\otimes\textbf{x}_{i}^{\top})\biggr\rvert\textbf{x}_{i}\right)\right)=0$
If the outcome model (9) is correctly specified, the third block in the first row of the Jacobian simplifies as follows:
$\mathbb{E}\left(-Z_{i}e^{-\textbf{x}_{i}^{\top}\bm{\gamma}_{0}}\bm{\epsilon}_{i}\textbf{x}_{i}^{\top}\right)=\mathbb{E}\bm{\epsilon}_{i}\mathbb{E}\left(-Z_{i}e^{-\textbf{x}_{i}^{\top}\bm{\gamma}_{0}}\textbf{x}_{i}^{\top}\right)=\textbf{0}.\\$
Next, calculate $\mathbb{E}[\psi_{i}(\bm{\theta}_{0})\psi_{i}(\bm{\theta}_{0})^{\top}]$. Still assuming both working models are correctly specified, we have:
$\mathbb{E}[\psi_{1,i}(\bm{\theta}_{0})\psi_{1,i}(\bm{\theta}_{0})^{\top}]$
$=\mathbb{C}\text{ov}\left[(\textbf{I}_{T\times T}\otimes\textbf{x}_{i}^{\top})\emph{vec}(\textbf{B}_{0})\right]+\mathbb{E}\left[\frac{Z_{i}}{\tau^{2}(\textbf{x}_{i}^{\top}\bm{\gamma}_{0})}\bm{\epsilon}_{i}\bm{\epsilon}_{i}^{\top}\right]$
$=\textbf{B}^{\top}_{0}\bm{\Sigma}_{x}\textbf{B}_{0}+\mathbb{E}[\tau^{-1}(\textbf{x}_{i}^{\top}\bm{\gamma}_{0})]\bm{\Sigma}_{\epsilon},$
@@
$\mathbb{E}\left[\frac{1}{\tau(\textbf{x}_{i}^{\top}\bm{\gamma}_{0})}\right]=\frac{1}{\pi}\left(1-\frac{\mathbb{V}(\tau(\textbf{x}_{i}^{\top}\bm{\gamma}_{0}))}{[\mathbb{E}\tau(\textbf{x}_{i}^{\top}\bm{\gamma}_{0})]^{2}}\right)=\frac{1}{\pi}\left(1-\frac{\mathbb{V}(\tau(\textbf{x}_{i}^{\top}\bm{\gamma}_{0}))}{\pi^{2}}\right),$
@@
$\mathbb{V}(\tau(\textbf{x}_{i}^{\top}\bm{\gamma}_{0}))$ $=\mathbb{E}_{\textbf{x}}[\tau(\textbf{x}_{i}^{\top}\bm{\gamma}_{0})^{2}]-\pi^{2}.$
@@
$\mathbb{E}[\psi_{1,i}(\bm{\theta}_{0})\psi_{2,i}(\bm{\theta}_{0})^{\top}]$
$=\mathbb{E}\left[\left((\textbf{I}_{T\times T}\otimes\textbf{x}_{i}^{\top})\emph{vec}(\textbf{B}_{0})-\bm{\mu}_{y}^{0}+\frac{Z_{i}}{\tau(\textbf{x}_{i}^{\top}\bm{\gamma}_{0})}\bm{\epsilon}_{i}\right)\left(\bm{\epsilon}_{i}\otimes Z_{i}\textbf{x}_{i}\right)^{\top}\right]$
$=\mathbb{E}\left[\frac{Z_{i}}{\tau(\textbf{x}_{i}^{\top}\bm{\gamma}_{0})}\bm{\epsilon}_{i}\left(\bm{\epsilon}_{i}\otimes Z_{i}\textbf{x}_{i}\right)^{\top}\right]=\mathbb{E}\left[\frac{Z_{i}}{\tau(\textbf{x}_{i}^{\top}\bm{\gamma}_{0})}(\bm{\epsilon}_{i}\bm{\epsilon}_{i}^{\top}\otimes\textbf{x}_{i}^{\top})\right]$
$=\bm{\Sigma}_{\epsilon}\otimes\bm{\mu}_{x}^{\top}$
@@
$\mathbb{E}[\psi_{1,i}(\bm{\theta}_{0})\psi_{3,i}(\bm{\theta}_{0})^{\top}]$
$=\mathbb{E}\left[\left((\textbf{I}_{T\times T}\otimes\textbf{x}_{i}^{\top})\emph{vec}(\textbf{B}_{0})-\bm{\mu}_{y}^{0}+\frac{Z_{i}}{\tau(\textbf{x}_{i}^{\top}\bm{\gamma})}\bm{\epsilon}_{i}\right)(Z_{i}-\tau(\textbf{x}_{i}^{\top}\bm{\gamma}))\textbf{x}_{i}^{\top}\right]=\textbf{0}.$
@@
$\mathbb{E}[\psi_{2,i}(\bm{\theta}_{0})\psi_{2,i}(\bm{\theta}_{0})^{\top}]$
$=\mathbb{E}\left[(\bm{\epsilon}_{i}\otimes Z_{i}\textbf{x}_{i})(\bm{\epsilon}_{i}\otimes Z_{i}\textbf{x}_{i})^{\top}\right]=\mathbb{E}\bm{\epsilon}_{i}\bm{\epsilon}_{i}^{\top}\otimes\left[Z_{i}\textbf{x}_{i}\textbf{x}_{i}^{\top}\right]$
$=\mathbb{E}\left[\bm{\epsilon}_{i}\bm{\epsilon}_{i}^{\top}\right]\otimes\mathbb{E}\left[Z_{i}\textbf{x}_{i}\textbf{x}_{i}^{\top}\right]=\bm{\Sigma}_{\epsilon}\otimes\bm{\Pi}$
@@
$\mathbb{E}[\psi_{2,i}(\bm{\theta}_{0})\psi_{3,i}(\bm{\theta}_{0})^{\top}]=\mathbb{E}\left[(\bm{\epsilon}_{i}\otimes Z_{i}\textbf{x}_{i})(Z_{i}-\tau(\textbf{x}_{i}^{\top}\bm{\gamma}))\textbf{x}_{i}^{\top}\right]=\textbf{0}.$
Since $\hat{\bm{\gamma}}$ is a maximum likelihood estimator, the blocks in $\mathbb{E}[\psi_{3,i}(\bm{\theta}_{0})\psi_{3,i}(\bm{\theta}_{0})^{\top}]$ are identical to the expected value of the corresponding blocks in the Jacobian. According to the M-estimation approach, we have $\sqrt{n}(\hat{\bm{\theta}}-\bm{\theta})\overset{d}{\rightarrow}\mathcal{N}(\textbf{0},\bm{\Sigma_{\theta}})$, where, by the sandwich estimator formula, the asymptotic variance of $\hat{\bm{\theta}}$ is
$\mathbb{E}\left(\frac{\partial\psi_{i}(\bm{\theta})}{\partial\bm{\theta}}\biggr\rvert_{\bm{\theta}_{0}}\right)^{-1}\mathbb{E}[\psi_{i}(\bm{\theta}_{0})\psi_{i}(\bm{\theta}_{0})^{\top}]\left\{\mathbb{E}\left(\frac{\partial\psi_{i}(\bm{\theta})}{\partial\bm{\theta}}\biggr\rvert_{\bm{\theta}_{0}}\right)^{-1}\right\}^{\top}$
$=\begin{bmatrix}\textbf{B}^{\top}_{0}\bm{\Sigma}_{x}\textbf{B}_{0}+\mathbb{E}[\tau^{-1}(\textbf{x}_{i}^{\top}\bm{\gamma}_{0})]\bm{\Sigma}_{\epsilon}&-\bm{\Sigma}_{\epsilon}\otimes\bm{\mu}_{x}^{\top}\bm{\Pi}^{-1}&\textbf{0}\\ \bm{\Sigma}_{\epsilon}\otimes\bm{\Pi}^{-1}\bm{\mu}_{x}&\bm{\Sigma}_{\epsilon}\otimes\bm{\Pi}^{-1}&\textbf{0}\\ \textbf{0}&\textbf{0}&\bm{\Phi}^{-1}\end{bmatrix}$.
@@
Lemma 4.
Assume the working model (1) is correctly specified, i.e., $E(\textbf{y}_{i}\mid\textbf{x}_{i})=\textbf{B}^{\top}\textbf{x}_{i}$. Then, the OR estimator (10) has the following asymptotic multivariate normal distribution:
$\sqrt{n}(\hat{\bm{\mu}}_{OR}-\bm{\mu}_{y})\overset{d}{\rightarrow}\mathcal{N}\left(\textbf{0},\textbf{B}^{\top}\bm{\Sigma}_{x}\textbf{B}+\bm{\mu}_{x}^{\top}\bm{\Pi}^{-1}\bm{\mu}_{x}\bm{\Sigma}_{\epsilon}\right),$
where $\bm{\Pi}=\mathbb{E}\left[Z_{i}\textbf{x}_{i}\textbf{x}_{i}^{\top}\right]$.
@@
Here, we apply the same approach to find the limiting distribution of outcome regression imputed estimator. The outcome regression imputed estimator is defined as $\hat{\bm{\mu}}_{OR}=n^{-1}\sum_{i=1}^{n}\hat{\textbf{B}}^{\top}\textbf{x}_{i}$. Let $\bm{\theta}=(\bm{\mu}_{y},\emph{vec}(\textbf{B}))^{\top}$. Define equations for calculating the estimator as:
$\psi_{1,i}(\bm{\theta})=(\textbf{I}_{T\times T}\otimes\textbf{x}_{i}^{\top})\emph{vec}(\textbf{B})-\bm{\mu}_{y},$ (15)
$\psi_{2,i}(\bm{\theta})=(\textbf{y}_{i}-(\textbf{I}_{T\times T}\otimes\textbf{x}_{i}^{\top})\emph{vec}(\textbf{B}))\otimes Z_{i}\textbf{x}_{i}.$ (16)
Calculate the Jacobian of $\psi_{i}(\bm{\theta})$ and evaluate it at the true value $\bm{\theta}_{0}$. The first row of Jacobian is
$\frac{\partial\psi_{1,i}(\bm{\theta})}{\partial\bm{\theta}^{\top}}\biggr\rvert_{\bm{\theta}_{0}}=\left[-\textbf{I}_{T\times T},\textbf{I}_{T\times T}\otimes\textbf{x}_{i}^{\top}\right].$
@@
$\frac{\partial\psi_{2,i}(\bm{\theta})}{\partial\bm{\theta}^{\top}}\biggr\rvert_{\bm{\theta}_{0}}=\left[\textbf{0}_{Tp\times T},-\textbf{I}_{T\times T}\otimes Z_{i}\textbf{x}_{i}\textbf{x}_{i}^{\top}\right].$
@@
$\mathbb{E}\left(\frac{\partial\psi_{i}(\bm{\theta})}{\partial\bm{\theta}}\biggr\rvert_{\bm{\theta}_{0}}\right)=\begin{bmatrix}-\textbf{I}_{T\times T}&\textbf{I}_{T\times T}\otimes\bm{\mu}_{x}^{\top}\\ \textbf{0}_{Tp\times T}&-\textbf{I}_{T\times T}\otimes\bm{\Pi}_{p\times p}\end{bmatrix}$,
where $\bm{\Pi}=\mathbb{E}\left[Z_{i}\textbf{x}_{i}\textbf{x}_{i}^{\top}\right]$. By the M-estimation approach, the asymptotic variance of $\hat{\bm{\theta}}$ is
$\mathbb{E}\left(\frac{\partial\psi_{i}(\bm{\theta})}{\partial\bm{\theta}}\biggr\rvert_{\bm{\theta}_{0}}\right)^{-1}\mathbb{E}[\psi_{i}(\bm{\theta}_{0})\psi_{i}(\bm{\theta}_{0})^{\top}]\left\{\mathbb{E}\left(\frac{\partial\psi_{i}(\bm{\theta})}{\partial\bm{\theta}}\biggr\rvert_{\bm{\theta}_{0}}\right)^{-1}\right\}^{\top}$
$=\begin{bmatrix}\textbf{B}^{\top}_{0}\bm{\Sigma}_{x}\textbf{B}_{0}+\bm{\mu}_{x}^{\top}\bm{\Pi}^{-1}\bm{\mu}_{x}\bm{\Pi}_{\epsilon}&\bm{\Sigma}_{\epsilon}\otimes\bm{\mu}^{\top}\bm{\Pi}^{-1}\\ \bm{\Sigma}_{\epsilon}\otimes\bm{\Pi}^{-1}\bm{\mu}_{x}&\bm{\Sigma}_{\epsilon}\otimes\bm{\Pi}^{-1}\end{bmatrix}$.
@@
Lemma 3 states the double robustness property of the DR estimator [e.g. 32]. However, up to our knowledge, the covariance matrices of the pointwise OR- and DR-estimators in the context of multivariate outcomes do not appear in the literature elsewhere. These covariance matrices are necessary to provide simultaneous inference as presented in Section 2.3.